How many gallons of a 60% antifreeze solution must be mixed with 100 gallons of 15% antifreeze to get a mixture that is 50% antifreeze?
Answer by stanbon(57984)

How many gallons of a 60% antifreeze solution must be mixed with 100 gallons of 15% antifreeze to get a mixture that is 50% antifreeze?
---------------------------
Amt of 60% solution = x gallons ; Amt of active ingredient = 0.6x gallons
-----------------------------------------------
Amt of 15% solution = 100 gallons ; Amt of active ingredient = 15 gallons
------------------------------------------------
Amt of 50% solution = 100+x gallons ; Amt of active = 50 + 0.5x gallons
------------------------------------------
EQUATION:
Active ingredient: 0.60x + 15 = 50 + 0.50x
0.10x = 35
x = 350 gallons (amt of 60% solution that must be added)
==========================
Cheers,
Stan H.